The two pendulums shown have 14 kg balls supported
Answer #1

given

m = 14 kg , angle \theta = 38o , R = 2 m

coefficient of restitution is = 0.6

we have initial velocity u and mass

then

m g ( R + R sin\theta ) = 1/2 m u2

u = ( 2 g R ( 1 + sin\theta ) )1/2

impulse can be defined as the change in momentum

so impulse = change in momentum

impulse = m ( u + u / 1.66 ) ( since used coefficient of restitution = 0.6 )

= 2.66 m u / 1.66

substituting the " u " value

impulse = 2.66 / 1.66 X m X ( 2 g R ( 1 + sin\theta ) )1/2   

= 2.66 / 1.66 X 14 X ( 2 X 9.8 X 2 X ( 1 + sin38 ) )1/2  

= 22.43 X 7.958

impulse = 178.527 kg.m/sec